# CoRaiS: Lightweight Real-Time Scheduler for Multi-Edge Cooperative Computing

### v1

@@ I Introduction
Many researchers are interested in the multi-edge cooperation system, and have proposed some algorithms [3, 4, 5, 6, 7, 8] to schedule independent requests and requests with logical execution order in the system. They usually make some hypothesises to support their research. Firstly, only CPU is configured on the edges, and the number of calculations required to respond to each request is known. With such two conditions, the response time of each request can be obtained by dividing the calculation number by the CPU frequency. Secondly, the request arrival patterns follow known probability distributions, such as poisson distribution, multinomial distribution, etc. With such assumption, many statistical theories and models can be applied to the multi-edge scheduling problems.
@@
Based on the analysis above, we can identify that the multi-edge cooperative scheduling faces three challenges. (i) Multi-edge cooperative computing system state modeling: The diverse hardware composition and different resource allocation schemes for services at edges pose a severe problem for modeling system state. The QoS of edges would not be fairly evaluated, unless the edges heterogeneity can be shielded and an unified QoS evaluation method can be built. Meanwhile, the system state that includes remaining resource of CPU/GPU/Memory at edges keeps changing as the arrival, running, and completion of requests. Dispatching requests to edges based on the perfect initial state, like traditional approaches, will make the scheduling solution deviate from the optimal at the current state. Therefore, it is challenging but necessary to build a system-level state model that supports edges unified modeling and perceives dynamic changes of resource states. (ii) Multi-edge cooperative scheduling formulation: Most previous theoretical models were designed based on probability distribution assumption of requests and single computing hardware assumption of edges [9, 10], which may not accurately reflect real-world application scenarios. Some learning-based approaches [4, 5] are also studied to optimize the scheduling behaviours on some specific datasets and specially constructed multi-edge network. However, the generalization ability [^2] Here the generalization ability refers to the pre-trained scheduling models can help make effective decisions on unseen applications and multi-edge networks. of the approaches is not good. Once the application and network environment change, a large amount of data (with expert knowledge sometimes) must be collected to retrain the scheduling model, while sometimes it is difficult to collect efficient data and expert knowledge. Therefore, a new mathematical formulation is required to reveal the essence of multi-edge cooperative scheduling problem and to guide scheduling algorithm research for requests with any arrival pattern. (iii) Real-time scheduler designing: The multi-edge scheduling problem for multiple requests is essentially a cominatorial optimization problem with some constraints. The search space of such problem is huge and will grow as the number of edges and requests increase. Computing an optimal solution in the huge search space is theoretically time-consuming. The previous works usually took a long time to make the scheduling decision. However, in practice, only methods that support real-time scheduling can ensure the efficient operation of the multi-edge cooperative computing system.
@@ II Related Work
Existing works have explored the formulation and optimization of scheduling problems in the multi-edge or edge-cloud cooperative computing system and gotten some great contributions [11, 12, 13]. They generally divide dispatched requests into two categories: independent requests and requests with logical execution order, and then design scheduling algorithms respectively.
The efforts [14, 3, 4, 5, 15, 9, 10, 16, 17] focus more on independent requests, as well as this paper. However, there are significant differences between these works and this paper in terms of modeling goal, problem settings and optimization approaches. The goal of works [14, 3, 4, 5] is to maximize the expected number of requests served per slot with different assumptions. The work [14, 3] assume that the computation requirement and average arrival rate of requests are known, formulates the problem as an integer linear program and prove the problem is generally NP-hard. A heuristic approximation algorithm based on linear program relaxation and rounding is proposed to solve the problem as well. The work [4] builds their own experiment environment that consists of some physical servers and leased clouds, and proposes a coordinated multi-agent actor-critic algorithm for decentralized request dispatch. The algorithm is evaluated on real-world workload traces from Alibaba. Another learning-based scheduling model EdgeMatrix is presented to maximize the throughput while guaranteeing various Service-Level-Agreement priorities in the work [5], where a networked multi-agent actor-critic algorithm to customize resource channels is proposed to improve the system’s stability. The work [15] formulates a request routing problem under the assumptions of computation capacity (i.e. maximum frequency) of a single CPU in the multi-cell mobile edge computing networks, and proposes a randomized rounding algorithm to minimizes the load to cloud. The work [9] assumes that the request arrival at each edge is a Poisson process, the computation requirements (in CPU cycle) of requests follow exponential distribution, and the cloud always has enough computation capacity to provide near-zero computing delay. The work [10] assumes the dynamic arrival of computation requests can be approximated as Poisson process, and the request requires the computing support by a random number of CPU cycles with a finite mean value. The authors transform the scheduling problem into a coopeartive queueing game approach to minimize the expected cost of each individual edge server. The works [16, 17] attach to the response time of each request a weight to indicate its latency sensitivity and introduce a machine-dependent processing time in each server. Then the works propose an online scheduling framework based on question and answer (Q&A) mode to minimize the total weighted response time over all requests.
There are some other progresses on dispatching requests with logical execution order [18, 19, 20, 6, 7]. The dependencies among such requests are usually formulated as directed acyclic graphs (DAGs), and the dependency-aware requests scheduling problems are also know as DAG scheduling problem. The work [18] proposes Decima that develops new representations for requests’ dependency graphs and adopts scalable reinforcement learning models to learn work-load-specific scheduling algorithms without any human instruction, in order to minimize average request completion time. The paper [19] allocates computing devices to continuous data flows in a large distributed system, through presenting a graph-aware encode-decoder framework to learn a generalizable resource allocation strategy. The works [20, 6, 7] present how to optimize device placement for training deep neural networks (DNN). The architecture of a DNN model can take as a DAG, and the device placement for DNN is to specify how each operation in a DNN model should be matched to the heterogeneous CPU and GPU devices. The [20] presents a sequence-to-sequence model that learns to optimize device placement for TensorFlow computational graphs. Spotlight is proposed [6] to find a optimal device placement for training DNNs. Spotlight models the problem as a Markov decision process with multiple stages, and uses a new reinforcement learning algorithm based on proximal policy optimization. A two-level hierarchical model [7] realizes device placement for a neural network with tens of thousands of operations.
In general, the scheduling problem can be taken as a combinatorial optimization problem under some constrains and is proved to be NP-hard. In recent years, designing algorithms to quickly solve combinatorial optimization problems has attracted much attention, and some achievements have been made. The works [21, 22] design learning-based model to solve traveling salesman problems (TSP). The advances [23, 24, 25] learn policies for vehicle routing problems (VRP). The works [26] propose reinforcement learning based models to solve multiple TSP. The progresses [27, 28] present common framework for combinatorial problems. Solutions for mixed integer programming (MIP) are proposed as well [29, 30].
III System Model and Problem Formulation
III-A Multi-edge Cooperative Scheduling Processes
Figure 2: Scheduling processes in multi-edge cooperative computing
The multi-edge cooperative computing system consists of a set of network edges =$\mathcal{E}=\{e_{n}\}_{n=1}^{N}$, and can provide diverse services $\mathcal{S}=\{s_{k}\}_{k=1}^{K}$. Each edge is under the exclusive management of a single central scheduler, effectively eliminating the possibility of ambiguous scheduling decisions. Furthermore, every edge within the system willingly collaborates with others and entrusts the central scheduler’s decision-making efficacy. Consequently, each edge actively accepts transmitted requests without rejection. This collective commitment ensures a seamless and efficient process, contributing to the overarching goal of achieving timely completion of all tasks. The processes of multi-edge cooperative scheduling are shown in Fig. 2. Before describing the detail processes, we first explain some concepts used in the scheduling process as follows.
Request: A request usually consists of description text and physical input data. The description text describes the required service, the client ID, etc. Input data is the data to be processed. For example, an image classification request should explain which classifier is required in the description text, and upload the related images as input data.
Request brief: It is a data package that only has some description information of request, without detailed input data. For example, the brief of an image classification request records the size of images and required classifier, and does not include the content of images. By introducing request brief, the data packages used to make scheduling decisions become very small, which would significantly reduce the communication delay among edges and central controller.
Scheduling decision: It contains the information which edge the requests would be executed at. The edges must obey the decision to locally respond to the requests or transfer the requests to other edges.
Then as shown in Fig. 2, the scheduling processes involve seven steps. (i) Clients submit requests and related data to edges. (ii) Edges receive requests and create a request brief for each of them. (iii) Edges submit current service capacity information and the briefs to central controller (CC). (iv) CC makes a scheduling decision based on edge status and request briefs, and designated scheduling algorithms. (v) CC informs edges about the decision, where the execution edge for each request is decided. (vi) Based on the decision, edges determine whether to locally respond to the requests or transmit them to other edges with the relevant data. (vii) The edges feed back the computing results to clients.
III-B System Scheduling Decomposition
@@
As the decomposition principle illustrated in Fig. 3, we decompose the multi-edge cooperative system into multiple service-oriented subsystem, i.e. $\{SR_{k}\}_{k=1}^{K}$, where $SR_{k}=\{e_{k1},e_{k2},...,e_{kn^{\prime}}\}$ and $e_{ki}$ denotes the $i^{th}$ edge with service $s_{k}$. With the decomposition, our focus can then shift towards addressing scheduling problem within each individual subsystem, since the process of scheduling over the whole multi-edge system can be decomposed into scheduling on each $SR_{k}$. For the sake of clarity in the following discussions, we omit explicitly declaring the $s_{k}$-oriented. However, it’s important to understand that all the operations described can be applied into any service-oriented subsystem.
III-C System-level State Evaluation Model

### v2

@@ I Introduction
Many researchers are interested in the multi-edge cooperation system, and have proposed some algorithms [3, 4, 5, 6, 7, 8, 9] to schedule independent requests and requests with logical execution order in the system. They usually make some hypothesises to support their research. Firstly, only CPU is configured on the edges, and the number of calculations required to respond to each request is known. With such two conditions, the response time of each request can be obtained by dividing the calculation number by the CPU frequency. Secondly, the request arrival patterns follow known probability distributions, such as poisson distribution, multinomial distribution, etc. With such assumption, many statistical theories and models can be applied to the multi-edge scheduling problems.
@@
Based on the analysis above, we can identify that the multi-edge cooperative scheduling faces three challenges. (i) Multi-edge cooperative computing system state modeling: The diverse hardware composition and different resource allocation schemes for services at edges pose a severe problem for modeling system state. The QoS of edges would not be fairly evaluated, unless the edges heterogeneity can be shielded and an unified QoS evaluation method can be built. Meanwhile, the system state that includes remaining resource of CPU/GPU/Memory at edges keeps changing as the arrival, running, and completion of requests. Dispatching requests to edges based on the perfect initial state, like traditional approaches, will make the scheduling solution deviate from the optimal at the current state. Therefore, it is challenging but necessary to build a system-level state model that supports edges unified modeling and perceives dynamic changes of resource states. (ii) Multi-edge cooperative scheduling formulation: Most previous theoretical models were designed based on probability distribution assumption of requests and single computing hardware assumption of edges [10, 11], which may not accurately reflect real-world application scenarios. Some learning-based approaches [5, 6] are also studied to optimize the scheduling behaviours on some specific datasets and specially constructed multi-edge network. However, the generalization ability [^2] Here the generalization ability refers to the pre-trained scheduling models can help make effective decisions on unseen applications and multi-edge networks. of the approaches is not good. Once the application and network environment change, a large amount of data (with expert knowledge sometimes) must be collected to retrain the scheduling model, while sometimes it is difficult to collect efficient data and expert knowledge. Therefore, a new mathematical formulation is required to reveal the essence of multi-edge cooperative scheduling problem and to guide scheduling algorithm research for requests with any arrival pattern. (iii) Real-time scheduler designing: The multi-edge scheduling problem for multiple requests is essentially a combinatorial optimization problem with some constraints. The search space of such problem is huge and will grow as the numbers of edges and requests increase. Computing an optimal solution in the huge search space is theoretically time-consuming. The previous works usually took a long time to make the scheduling decision. However, in practice, only methods that support real-time scheduling can ensure the efficient operation of the multi-edge cooperative computing system.
@@
The following paper is structured as follows: Section II presents an overview of related work. In Section III, we provide two use cases to illustrate the key aspects of request scheduling in a multi-edge cooperative computing system. Sections IV and V delve into the designed model of the multi-edge cooperative computing system and the formulation of multi-edge scheduling, respectively. Following this, Section VI elaborates on the architecture of the proposed CoRaiS system. Simulation experiments and prototype-based experiments are detailed in Sections VII and VIII, respectively. Finally, the conclusion in Section IX summarizes the findings and contributions of the paper.
@@ II Related Work
Previous studies have delved into the formulation and optimization of scheduling problems within multi-edge or edge-cloud cooperative computing systems, yielding significant contributions [12, 13, 14]. Typically, these works categorize dispatched requests into two groups: independent requests and requests with a logical execution order, subsequently designing specific scheduling algorithms for each category.
II-A Independent Requests Scheduling
Numerous efforts [15, 4, 5, 6, 16, 10, 11, 17, 18] have primarily focused on independent requests. Most of them formulate the request scheduling problem within multi-edge cooperative system as integer linear programming problem, and make some assumptions to enable the formulations. For instance, [15, 4] assume known computation requirements and average arrival rates, formulating the problem as an integer linear program and proving its NP-hardness. They further propose heuristic approximation algorithms based on linear program relaxation and rounding to address the issue. Other works like [10, 11] make assumptions about request arrival patterns and computation requirements to address scheduling challenges. The works [17, 18] assume weightings indicating latency sensitivity and machine-dependent processing times for requests and propose online scheduling frameworks to minimize total weighted response time across all requests. The work [16] formulates a request routing problem under the assumption of computation capacity in multi-cell mobile edge computing networks, proposing a randomized rounding algorithm to minimize load to the cloud.
Besides the heuristic methods, learning-based approaches have emerged as a significant contributor to addressing multi-edge cooperative computing challenges. For instance, the work [5] develops its dedicated experimental environment comprising physical servers and leased clouds, introducing a coordinated multi-agent actor-critic algorithm for decentralized request dispatch. This algorithm is evaluated using real-world workload traces from Alibaba. Similarly, EdgeMatrix [6] presents a learning-based scheduling model aiming to maximize throughput while ensuring various Service-Level-Agreement priorities. A networked multi-agent actor-critic algorithm is proposed to customize resource channels and enhance system stability. The work [19] proposes a task scheduling framework across edge clouds, namely LsiA3CS, which employs A3C [20] and heuristic guidance to achieve distributed, asynchronous task scheduling for large-scale IIoT. Aiming at optimizing the average response time of task scheduling and the average energy consumption of the system, a multi-objective task scheduling model is designed, and a task scheduling policy optimization algorithm based on improved asynchronous advantage actor-critic (A3C) is proposed in [21].
Additionally, the optimization objective of researches are also varying. The work [22] proposes a fine-grained two-sided market via an intercloud broker, named SkyPilot to allow users to view the cloud ecosystem and schedule requests in an economic approach. The work [23] uses mixed integer programming to formulate the scheduling problem, which aims to satisfy the maximum number of requests given their deadline requirements. Moreover, the work presents a heuristic approach using the genetic algorithm to generate the scheduling decision. In the work [24], an optimal control theory (OCTS) based method is introduced. It aims to achieve global optimal results, ensuring computational and transmission latency requirements are met while load balancing various vehicle tasks. The work [25] proposes an evolutionary genetic (GA)-based optimization algorithm, aiming to strike a balance between task execution time and processing cost. The work [26] presents a Dynamic Parallel Computing Offload and Energy Management (DPCOEM) algorithm. It utilizes Lyapunov optimization technology to minimize task response time and achieve near-optimal performance. The work [27] formulates a bilevel optimization problem and proposes ant colony system and monotonic optimization method to minimize the energy consumption of all requests with deadline. The work [28] introduces a resource-constrained task scheduling profit optimization algorithm (RCTSPO). It focuses on profit maximization while enhancing task scheduling time, reliability, and load balancing.
TABLE I: Synthesis table of related works
Related works Computing resources Assumptions of request arrival pattern Assumptions of execution time for each request Problem formulations Optimization objectives Scheduling algorithms
[15, 4] CPU The average arrival rate at time slots is known N/A Mixed integer linear program (MILP) Maximize the number of requests served per slot Greedy algorithm and linear program
[10] CPU The arrival pattern at each edge node follows a Poisson process Requests’ CPU cycles follow exponential distribution. The computing time is estimated via FDC, i.e. the maximum frequency of edges divides required CPU cycles. Mixed integer non-linear programming problem (MINonILP) Minimize the service response time and outsourcing traffic to central clouds Iterative optimization algorithm
[11] CPU The average arrival rate at time slots is known Requests’ CPU cycles are randomly sampled with a finite expection. The computing time is estimated via FDC. Queuing game Minimize the expected cost of each individual edge server Cooperative queueing game approach
[17, 18] CPU N/A Requests’ CPU cycles are known and an edge can execute at most one request at a time Dual integer programming Minimize the total weighted response time of all requests Online algorithm with dispatching policy and the scheduling policy
[23] CPU N/A Requests’ CPU cycles are known, and the computing time is estimated via FDC. MILP Maximize the number of requests with deadlines Genetic algorithm
[26] CPU The arriving probability at each slot is known Requests’ CPU cycles are known, and the computing time is estimated via FDC. MINonILP Minimize response time and packet losses of tasks under the limitation of energy queue stability Lyapunov-based dynamic parallel computing offloading
[25] CPU N/A Requests’ CPU cycles are known, and the computing time is estimated via FDC. MILP Achieve a trade-off between execution time and monetary cost to complete requests Particle swarm optimization
[27] CPU N/A Requests’ CPU cycles are known, and the computing time is estimated via FDC. MILP Minimize the total energy consumption of all mobile users under the delay constraint Ant colony and monotonic optimization
[28] CPU N/A Requests’ CPU cycles are known, and the computing time is estimated via FDC. Petri Net Load balancing and profit optimization Heuristic algorithm
[5] CPU Specified based on dataset Obtained from dataset Markov game formulation Balance the workloads among edge and offload some requests to the cloud Coordinated multi-agent actor-critic
[6] CPU Specified based on dataset Obtained from dataset Markov game formulation and MILP Maximize the throughput while guaranteeing various SLA priorities Multi-agent actor-critic algorithm
[19] CPU Specified based on dataset Obtained from dataset Markov game formulation Maximize the long-term throughput rate Actor-critic algorithm
[29] CPU Stochastic requests arrivals The average execution time is known (from dataset) Markov decision process (MDP) Minimize average request completion time Policy gradient
[30] CPU N/A Requests’ CPU cycles are known, and the computing time is estimated via FDC. Search problem Maximize system throughput Graph-aware encoder-edecoder reinforcement learning model
[31] CPU N/A Requests’ CPU cycles are known, and the computing time is estimated via FDC. MDP Minimize the long-term average delay and energy consumption, and maximize the system throughput Value decomposition multi-agent deep Q learning
[32, 7, 8] GPU and CPU N/A Measured in the hardware environment MDP Minimize the training time of a deep neural network Reinforcement learning (RL)
[33] GPU and CPU N/A Measured in the hardware environment MILP and MDP Minimize the training time of deep neural networks Multi-agent RL
Ours Service-oriented any resource combinations N/A Estimated through the proposed system-level state evaluation model MILP Minimize the average response time of computing requests RL
II-B Scheduling Requests with Logical Execution Order
Dispatching requests with logical execution orders constitutes an important area of study within multi-edge cooperative computing. The dependencies among such requests are often represented as directed acyclic graphs (DAGs), leading to what is commonly referred to as the DAG scheduling problem.
The work [29] introduces Decima, which devises novel representations for requests’ dependency graphs. It employs scalable reinforcement learning models to learn workload-specific scheduling algorithms autonomously, aiming to minimize average request completion time. In [30], computing devices are allocated to continuous data flows within a large distributed system. The paper presents a graph-aware encode-decoder framework designed to learn a generalizable resource allocation strategy. The work [31] proposes an online concurrent user request scheduling mechanism based on multi-agent deep reinforcement learning to optimize the long-term average delay and energy consumption.
In addition to traditional request scheduling, the optimization of device placement for decomposed deep neural networks (DNNs) has also garnered attentions in the multi-edge cooperative scheduling domain, inspired by advancements in neural network-based intelligent large models. Studies also treat the architecture of a DNN model as a DAG, with device placement specifying how each operation in the model should be matched to networked heterogeneous CPU and GPU devices. For example, Mirhoseini et al. [32] propose a sequence-to-sequence model that optimizes device placement for TensorFlow computational graphs. Spotlight [7] is devised to determine optimal device placement for DNN training. It formulates the problem as a Markov decision process with multiple stages and employs a novel reinforcement learning algorithm based on proximal policy optimization. A two-level hierarchical model [8] is also introduced for device placement in neural networks containing tens of thousands of operations. The work [33] proposes TapFinger, a distributed scheduler for edge clusters that minimizes the total completion time of ML tasks through co-optimizing task placement and fine-grained multi-resource allocation.
II-C Related Work Analysis
We analyze the related works from perspectives of computing resources, assumptions on request arrival patterns and computation requirements, problem formulations, optimization objectives and scheduling algorithms, and summarized them in Table I. Delving into the details of previous researches, we can find that previous research often made assumptions based on simplified scenarios, such as only having CPUs on the edges and knowing the number of calculations required for each request. However, in reality, heterogeneous edges equipped with various computing units, including GPUs and TPUs, are becoming increasingly common. Moreover, with the rise in requests related to deep learning, deploying GPUs at the edges has emerged as a significant trend. This shift introduces a new layer of complexity. The interaction of multiple computing resources (CPU, GPU, TPU, etc.), each with its own capabilities and characteristics, coupled with the intricate execution logic of intelligent applications, poses challenges in accurately estimating the running time of requests. Additionally, many works in the field make assumptions about request arrival patterns following known probability distributions like Poisson or multinomial distributions. However, the reality is that the arrival pattern of requests is often unpredictable. Each client may indeed have its own unique request generation pattern, making it challenging to analyze the composite pattern of multiple clients.
In contrast to existing works, our research liberates from many assumptions. Firstly, we extend computing devices to encompass a combination of CPUs and GPUs, facilitating the development of complex intelligent applications. We achieve this by constructing a novel system-level state evaluation model capable of evaluating the computation time of each request. This model is compatible with heterogeneous hardware configurations and diverse intelligent applications. Secondly, we do not constrain the request arrival pattern. Instead, we design a new multi-edge cooperative framework and propose a corresponding lightweight scheduling algorithm to enable real-time scheduling of requests. Both the framework and algorithm are designed to accommodate any arrival pattern of requests.
III Motivating Use Cases
Figure 2: An illustration of the motivating use case.
Figure 3: An illustration of the motivating use case with mobility.
We present two use cases to illustrate the key points of request scheduling in multi-edge cooperative computing system. As shown in Fig. 2, five edges coordinate to serve clients, but the service capability and system status of edges are different, i.e., the hardware and software configure may be heterogeneous, and the backlogs and deployed service replicas can be different as well. Meanwhile, clients have different number of requests that need to be processed. Fig. 3 considers client mobility into the use case, i.e. clients can move to another client region after submitting their requests. Request scheduling has to dispatch each request to an edge and make the total response time of requests minimized with consideration of the edge heterogeneity and client mobility.
Multi-edge cooperative computing can also offer significant benefits in the context of content delivery networks (CDNs) [34, 35, 36], which are distributed networks of servers strategically positioned worldwide to deliver web pages, images, videos, and digital assets with high performance and availability. For instance, in video streaming applications, traditional CDN setups involve caching video content at edge servers for faster delivery. However, video content often requires transcoding into different formats or bitrates to suit various devices and network conditions. Multi-edge computing enables real-time video transcoding at edge servers. When a user requests a video, the edge server dynamically transcodes it into the appropriate format or bitrate based on the user’s device and network conditions. This eliminates the demands for centralized transcoding servers and reduces latency by delivering optimized video directly from the edge server. Similarly, in online gaming, CDNs optimize content delivery by caching game assets, updates, and patches at edge servers. Multi-edge computing can enhance this by distributing these assets across multiple edge locations, ensuring faster downloads and updates for players. Moreover, edge servers can dynamically adjust content delivery based on players’ location, network conditions, and device capabilities, ensuring optimal performance for each player.
However, there are two challenges when realizing the real-time effective scheduling. Firstly, there is a lack of effective business models or interoperability models to facilitate the cooperative interaction among edges and the precise execution of the scheduling decision. Because the traditional cloud business pattern, which submits all requests to the cloud management platform and lets the platform schedule the requests to specific servers, suffers from the huge data traffic consumption, the platform bottleneck and longer response time. Collaboration based on distributed consensus can lead to long-term negotiations at the edges, resulting in delayed scheduling decisions. Secondly, the scheduling space is $N^{M}$, where $N$ and $M$ refer to the number of edges and requests, respectively, indicating that the solution space will increase exponentially with the growth of edge numbers and request numbers. The huge scheduling space makes it difficult to search the optimized solution in time, while the practical scenario requires immediate solutions to enable the multi-edge system operation.
In the following, we will describe our interoperability models for the multi-edge cooperative computing system and formulate the scheduling problems in detail.
IV Multi-edge Cooperative Computing System Model
IV-A Service-oriented Multi-edge Computing System Modeling
@@
The multi-edge cooperative computing system consists of a set of network edges =$\mathcal{E}=\{e_{n}\}_{n=1}^{N}$, and can provide diverse services $\mathcal{S}=\{s_{k}\}_{k=1}^{K}$, as shown in Fig. 4. To enable service-oriented effective scheduling, we decompose the multi-edge cooperative system into multiple service-oriented subsystems, i.e. $\{SR_{k}\}_{k=1}^{K}$, where $SR_{k}=\{e_{k1},e_{k2},...,e_{kn^{\prime}}\}$, $e_{ki}$ denotes the $i^{th}$ edge with service $s_{k}$. With the decomposition, our focus can then shift towards addressing scheduling problem within each individual subsystem, since the process of scheduling over the whole multi-edge system can be decomposed into scheduling on each $SR_{k}$. For the sake of clarity, in the following discussions, we omit explicitly declaring the $s_{k}$-oriented. However, it’s important to understand that all the described operations can be applied into any service-oriented subsystems.
Figure 5: Territorial schedulers are introduced into the service-oriented multi-edge cooperative computing system to facilitate effective and real-time scheduling.
We further introduce territorial schedulers into the service-oriented multi-edge cooperative computing system. The novel multi-edge cooperative architecture is presented in Fig. 5. All edges are networked, which enabling data transmission between any two edges. Each edge is under the exclusive management of a single scheduler, which decides the execution location of each received request in its control region. Moreover, every edge within the system willingly collaborates with others and entrusts its scheduler’s decision-making efficacy. Consequently, each edge actively processes all requests dispatched by the scheduler. This collective commitment ensures a seamless and efficient process, contributing to the overarching goal of achieving timely completion of all tasks.
In addition, when an edge device seeks to join the system and contribute its computational resources, it must actively establish a connection with a base station. Subsequently, it is required to transmit registration messages to the territorial scheduler. These registration messages should include details such as deployed services, the number of corresponding replicas, computational capabilities, and geographical locations. Following the registration process, the edge device is expected to periodically send its status information to its designated scheduler, indicating its availability to handle incoming requests. If the information fails to be received continuously for a predetermined threshold number of times, the scheduler will infer that the edge device has gone offline, then if the edge wants to join the system again, it has to register like a new edge. By performing these operations, clients have the capability to submit requests to the edge device through either wireless or wired communication channels. Simultaneously, the scheduler is empowered to efficiently dispatch requests to the appropriate edge devices
IV-B Interoperability Model of Multi-edge Cooperative Scheduling
IV-B1 Concept definition
Some concepts used in the interoperability model of multi-edge cooperative scheduling are described as follows.
Request: A request typically comprises description text and physical input data. The description text outlines essential details such as the required service, client ID, source edge location from which the request originated, and the predicted edge location where the client will be located upon request completion. Input data refers to the actual data intended for processing. For instance, in an image classification request, the description text would specify the required classifier, while the related images would be uploaded as input data
Request brief: A request brief is essentially a data package containing only descriptive information about the request, devoid of detailed input data. For instance, in the context of an image classification request, the brief would encompass details such as the image sizes and the required classifier, while omitting the actual image content. Additionally, the brief includes information about the source edge location and the predicted edge location. The introduction of request briefs leads to a significant reduction in the size of data packages used for scheduling decisions. This reduction in package size has a direct impact on communication delay among edges and the scheduler, resulting in improved efficiency.
Scheduling decision: It contains the information which edge will execute the requests. The edges are required to adhere to this decision, either responding to the requests locally or transferring them to other edges as directed.
IV-B2 Interoperability scheduling model
Based on the architecture shown in Fig. 5, we further propose an interoperability scheduling model to enable the effective multi-edge cooperation.
The multi-edge cooperative scheduling processes involve eight steps: (i) Clients submit requests and related data to edges, along with their predicted edge location after request completion. (ii) Edges receive requests and generate a request brief for each. (iii) Edges provide their schedulers with current service capacity information and the request briefs. (iv) The scheduler makes scheduling decisions based on edge status, request briefs, and designated scheduling algorithms. (v) The scheduler informs edges of the decision, specifying the execution edge for each request. (vi) Based on the decision, edges locally handle the requests or transfer them to other edges along with the relevant data. (vii) The execution edges provide computing results feedback to the predicted edges. (viii) The predicted edges transmit the computing results to clients.
Figure 6: An example of multi-edge cooperative scheduling considering client mobility within a single control region. In Scheduler A’s control region, the submitted request is executed locally, indicating the source edge and the execution edge are the same. In Scheduler B’s control region, the request is transferred to another edge for processing. In this scenario, Scheduler B must simultaneously inform both the source edge and the execution edge (actions 5.1 and 5.2). Subsequently, the source edge transmits all relevant data of the request to the execution edge (action 6.1), which then follows the scheduling decision and handles the request accordingly (action 6.2). Once the request is completed, the execution edge transmits the results to the prediction edge (action 7), which in turn provides feedback to the client (action 8).
Figure 7: In a scenario involving multi-edge cooperative scheduling and client mobility across multiple control regions, schedulers are limited to scheduling edges within their respective control regions. When a request extends beyond this boundary, the scheduler dispatches it to one of its managed edges. Upon completion of the request, the execution edge transmits the results to the predicted edge via cross-region communications (action 7).
To provide a clearer illustration of the interaction processes between schedulers and edges supporting effective request response, we present two examples shown in Fig. 6 and Fig. 7. These examples encompass scenarios involving client mobility both within the control region of a scheduler and beyond its boundaries.
V Multi-edge Scheduling Formulation
In this section, we will formulate the multi-edge scheduling problem in detail.
V-A System-level State Evaluation Model